OpenStudy (anonymous):

Anyone good with pre calculus honors?? I have a test tomorrow on limits and I have to justify every answer! I'm confused

OpenStudy (anonymous):

have you done the proofs using epsilon/delta notation?

OpenStudy (anonymous):

No!!

OpenStudy (anonymous):

ok then they wont be on the test lol. you know the limit laws yes?

OpenStudy (anonymous):

The test is more like Lim. 1- square root of x/ x-1 X---> 1

OpenStudy (anonymous):

I can't figure out how to do all the steps, the answer is -1/2

OpenStudy (anonymous):

\[f(x)=x^2+4x\\ \lim_{x \rightarrow 2}f(x)=\lim_{x \rightarrow 2}(x^2+4x)\\=\lim_{x \rightarrow 2}x^2+\lim_{x \rightarrow2}4x\\=(\lim_{x \rightarrow 2}x)^2+4(\lim_{x \rightarrow 2}x)\\=(2)^2+4(2)\\=12\]

OpenStudy (anonymous):

do you understand that?

OpenStudy (anonymous):

Yah you can just plug the 2 in and you get 12 but all mine are x--> 0 and if you plug 0 in you get 0/0 so we have to factor cancel and other stuff!! How do I send a photo?

OpenStudy (anonymous):

oh so you know all the limit laws then, that's good. and just click "attack file" just below this the text box"

OpenStudy (anonymous):

and in your example, do you mean \[\lim_{x \rightarrow 1}(1-\sqrt{\frac{x}{x-1}})\] or \[\lim_{x \rightarrow 1}(1-\frac{\sqrt{x}}{x-1})\]?

zepdrix (zepdrix):

I think you meant, \[\large \lim_{x \rightarrow 1}\; \frac{1-\sqrt x}{x-1}\] like that, yes?

OpenStudy (anonymous):

@Amandamorris Which one from above do you mean?

OpenStudy (anonymous):

zep, I like that one a lot better lol

OpenStudy (anonymous):

The one for zep.. The last one

OpenStudy (anonymous):

I know you have to multiple by 1+ square root of x but where does the canceling begin?

OpenStudy (anonymous):

well multiply by the conjugate, (the \[\frac{1+\sqrt{x}}{1+\sqrt{x}}\]

OpenStudy (anonymous):

then

OpenStudy (anonymous):

you get a 1-x on the top and x-1 on the bottom, factor out a -1 from the top and now you have (-(x-1))/(x-1) that you can cacel

OpenStudy (anonymous):

Ok.. The answer is -1/2 and I'm just not getting that

OpenStudy (anonymous):

so \[\lim_{x \rightarrow 1}\frac{1-\sqrt{x}}{x-1}\\=\lim_{x \rightarrow 1}\frac{1-\sqrt{x}}{x-1}(\frac{1+\sqrt{x}}{1+\sqrt{x}})\\=\lim_{z \rightarrow 1}\frac{1-x}{(x-1)(1+\sqrt{x})}\\=\lim_{x \rightarrow 1}\frac{-(x-1)}{(x-1)(1+\sqrt{x})}\\=\lim_{x \rightarrow 1}\frac{-1}{1+\sqrt{x}}\\frac{-1}{2}\]

OpenStudy (anonymous):

the last line in my above post should be: \[=\frac{-1}{2}\]

OpenStudy (anonymous):

Oh I get it!! Thank you!!

OpenStudy (anonymous):

no problem :) sorry it took so long, I just hate not typing up equations in that form with latex

OpenStudy (anonymous):

Are you a high school student

OpenStudy (anonymous):

nope. :)

OpenStudy (anonymous):

How old r u

OpenStudy (anonymous):

21

OpenStudy (anonymous):

College? What college??

OpenStudy (anonymous):

im going to be going to either UWO or waterloo.

OpenStudy (anonymous):

Nice! Well thanks for the help!!

OpenStudy (anonymous):

no problem!